# Decomposing Uncertainty in Probabilistic Knowledge Graph Embeddings: Why Entity Variance Is Not Enough

### v1

@@ 1 Introduction
Knowledge graph (KG) embeddings enable efficient reasoning over structured knowledge (bordes2013translating; trouillon2016complex). As KGs evolve—new entities emerge, existing entities appear in novel relationships—robust uncertainty quantification becomes critical.
@@
Recent probabilistic methods learn distributions over embeddings (chen2019embedding; chen2021probabilistic; Chen2021PERM), where rare entities have higher variance. But these variances are relation-agnostic: entity $e$ receives a single variance $\sigma^{2}_{e}$ regardless of relational context. Consider $e$ appearing in 1000 triples with relations $\{r_{1},\ldots,r_{10}\}$ but never with $r_{11}$. The model assigns low variance (well-observed), yet queries $(?,r_{11},e)$ involve genuinely uncertain predictions.
This causes systematic failures: score-based methods achieve 0.99 AUROC detecting random corruptions (implausible triples) but only 0.52–0.64 on temporal distribution shift (queries from future time periods with new entities and relationships).
@@
Our contributions are threefold: (1) Theoretical: We prove that relation-agnostic uncertainty estimators achieve near-random performance on novel contexts (Theorem 1), and formalize conditions under which semantic and structural uncertainties are complementary (Theorem 2). (2) Empirical: We demonstrate this limitation persists across existing probabilistic methods (GP-KGE, UKGE, ensembles), achieving only 0.52–0.64 AUROC on temporal shift despite 0.99 on random corruptions. (3) Methodological: We propose CAGP (Coverage-Augmented GP-KGE), which augments entity-level variance with explicit coverage tracking of entity-relation co-occurrence. CAGP achieves 0.94–0.99 AUROC on temporal OOD across multiple benchmarks, a 60–80% relative improvement over probabilistic baselines.
@@ 2 Related Work
UKGE (chen2019embedding) associates confidence scores with triples; BEUrRE (chen2021probabilistic) uses box embeddings where volume indicates uncertainty; GP-KGE (Chen2021PERM) learns Gaussian distributions over entities. All share the limitation we identify: learned variances are relation-agnostic. Despite being trained on data containing entity-relation co-occurrence patterns, these methods do not autonomously learn relation-specific uncertainty, necessitating our explicit decomposition framework.
@@
MC Dropout (gal2016dropout), Deep Ensembles (lakshminarayanan2017simple), and Energy-based methods (liu2020energy) capture model uncertainty but not structural observation patterns. SNGP (liu2020simple) estimates distance to training data but remains relation-agnostic. Graph Posterior Networks (stadler2021graph) propagate uncertainty through graphs but treat edges homogeneously.
@@ 3 Background
Given a trained model, distinguish in-distribution (ID) test triples from out-of-distribution (OOD) corruptions. The standard protocol (safavi2020codex) generates OOD triples by replacing the tail with a random entity: $(h,r,t)\to(h,r,t^{\prime})$ where $t^{\prime}\sim\text{Uniform}(\mathcal{E})$. However, this creates easily detectable corruptions (low model scores). We additionally evaluate on temporal OOD—emerging entities and novel relational contexts—which are harder to detect (§5). Performance is measured by AUROC.
@@
Probabilistic embedding methods (Chen2021PERM) learn Gaussian distributions over entity embeddings:
@@ 4.1 Preliminaries
We set $\tau$ to the 10th percentile of entity frequencies (robust to outliers; see Appendix I for sensitivity analysis).
@@ 4.3 Theoretical Results
Under assumptions A1–A3 (variance-frequency monotonicity, ID coverage completeness, frequency overlap; see Appendix [^1] We empirically validate Assumption A3 in Appendix G.1, finding 100% of novel-context triples have frequency-matched ID counterparts on three benchmarks.), there exists a distribution $\mathcal{D}$ with novel-context OOD subset $\mathcal{D}_{\text{novel}}$ such that:
@@
Under idealized conditions (monotonic variance-frequency relationship, complete ID coverage, approximate frequency overlap; see Appendix A for precise statements and robustness analysis):
@@
Proof sketch. Part (i): Novel-context entities have high frequency (by definition), so their variance matches ID entities—semantic uncertainty cannot distinguish them. Part (ii): Emerging entities are defined by low frequency, but may still have coverage for the specific query relation, making some indistinguishable from ID via coverage alone. Part (iii): Novel contexts have missing coverage by definition, while ID has full coverage. Part (iv): The combination uses semantic to catch emerging entities that structural misses (those with coverage), and structural to catch novel contexts that semantic misses. Full proof in Appendix A.
4.4 Coverage-Augmented GP-KGE (CAGP)
@@ 5.1 Experimental Setup
We evaluate on three benchmarks with temporal-like OOD splits: FB15k-237 (14,541 entities, 237 relations), WN18RR (40,943 entities, 11 relations), and YAGO3-10 (123,161 entities, 37 relations). We create temporal-like distribution shifts by partitioning test sets based on entity frequency and entity-relation co-occurrence patterns, simulating the key temporal challenge: detecting when familiar entities appear in unfamiliar relational contexts (methodology detailed in Appendix H).
@@
Score-based (UKGE (chen2019embedding), Energy (liu2020energy)), ensemble (MC Dropout (gal2016dropout), Deep Ensemble (lakshminarayanan2017simple)), distance-aware (SNGP (liu2020simple)), and single signals ($U_{\text{sem}}$, $U_{\text{str}}$).
@@ 5.4 Standard OOD Detection
Theorem 2 predicts semantic AUROC =0.5=0.5 on novel contexts, but Table 2 shows 0.42. This discrepancy arises because assumption (A3) (frequency overlap) is approximate: novel-context entities tend toward higher frequencies than the ID average, causing slightly lower semantic uncertainty than ID. The prediction remains directionally correct (near-random).
@@
To characterize CAGP’s failure modes, we analyzed 40,000 predictions on FB15k-237 standard OOD (random tail corruption: 20,000 ID + 20,000 corrupted). The model achieves 0.971 AUROC with 91.3% accuracy. False positives (1,733, 8.7% of ID) occur primarily on triples with low-degree tail entities (mean degree 183 vs 583 for correctly classified ID) and rare relations (mean frequency 2,384 vs 5,213). False negatives (1,733, 8.7% of OOD) involve corrupted tails that coincidentally create higher-degree entities than typical corruptions (mean degree 183 vs 37). The balanced error rates (8.7% FP = FN) and high AUROC validate that both uncertainty components contribute meaningfully. See Appendix F for detailed breakdown.
See Appendix B for additional ablations and Appendix C for calibration analysis.
@@ 6 Conclusion
CAGP’s coverage matrix $\mathbf{C}\in\{0,1\}^{|\mathcal{E}|\times|\mathcal{R}|}$ requires $O(|\mathcal{E}|\times|\mathcal{R}|)$ memory. For our evaluated datasets (FB15k-237: 13MB dense, <<1MB sparse; YAGO3-10: 17.5MB dense, <<1MB sparse), this is negligible. For web-scale KGs, sparse storage (hash tables storing only observed pairs) reduces memory to $O(|\mathcal{T}|)$ with $O(1)$ inference. We also explored learned relation-conditioned variance via MLP, but found explicit coverage simpler and equally effective. CAGP provides interpretability and simplicity suitable for domain-specific KGs with <<1M entities; empirical evaluation on billion-triple KGs remains future work.
@@
Appendix A Theorem Proofs
@@
A.1 Proof of Theorem 1 (Impossibility Result)
@@
A.2 Proof of Theorem 2 (Complementarity)
@@
The assumptions (A1)–(A6) provide idealized conditions ensuring sharp performance guarantees. In practice, these assumptions may be violated to varying degrees. Table 12 (Appendix G) quantifies violations across datasets: (A1) Spearman correlation ranges from 0.74-0.74 to 0.85-0.85 (not perfect monotonicity); (A4) bounded semantic gap $\Delta$ exceeds 1.0 on WN18RR ($\Delta=1.10$) and YAGO ($\Delta=1.01$).
@@
Appendix B Ablation Study
Table 5: Ablation of mixing strategies on FB15k-237.
Configuration Temporal OOD Standard OOD
Fixed $\alpha=0.5$ 0.951 0.935
Learned global $\alpha$ (CAGP) 0.986 0.960
B.1 Binary vs Continuous Coverage
Table 6: Binary vs continuous coverage on temporal OOD (FB15k-237). Binary achieves perfect detection while continuous variants fail.
Coverage Mode AUROC AUPR Separation
Binary 1.0000 1.0000 0.360
Log-scaled 0.5888 0.6119 0.058
TF-IDF 0.5606 0.5733 0.046
Explanation.
Why perfect detection on some temporal splits?
Binary coverage achieves perfect AUROC =1.0=1.0 on FB15k-237’s simulated temporal split (Table 6 above) but only 0.824 on ICEWS14 (main text Table 1). This discrepancy reflects different temporal split compositions:
Dataset Novel Contexts Emerging Entities
FB15k-237 (simulated) $\sim$94% $\sim$6%
ICEWS14 (ground-truth) $\sim$61% $\sim$39%
Appendix C Calibration Analysis
Method ECE $\downarrow$ Brier $\downarrow$
UKGE 0.312 0.298
Energy 0.287 0.271
Deep Ensemble 0.198 0.203
SNGP 0.183 0.195
$U_{\text{sem}}$ 0.156 0.178
$U_{\text{str}}$ 0.142 0.165
CAGP 0.089 0.112
Appendix D Selective Prediction
Method Acc@85% Error Red.
No abstention 62.3%
UKGE 68.1% 15.4%
Deep Ensemble 69.4% 18.8%
$U_{\text{sem}}$ 71.2% 23.6%
$U_{\text{str}}$ 73.5% 29.7%
CAGP 78.4% 42.7%
Appendix E Architecture Generalization
Table 10: Generalization across architectures on FB15k-237.
Base Model $U_{\text{sem}}$ $U_{\text{str}}$ CAGP
DistMult 0.753 0.821 0.959
TransE 0.775 0.822 0.963
ComplEx 0.755 0.821 0.960
Appendix F Error Analysis
Metric Value Rate
Overall Performance
AUROC 0.971
Accuracy 91.3%
Precision 0.913
Recall 0.913
F1 Score 0.913
Confusion Matrix
True Negatives 18,267
False Positives 1,733 8.7%
False Negatives 1,733 8.7%
True Positives 18,267
False Positive Characteristics
Avg tail degree 183 (ID: 583)
Avg relation freq 2,384 (ID: 5,213)
False Negative Characteristics
Avg tail degree 183 (OOD: 37)
Key findings.
False positives (8.7%) occur primarily on ID triples with low-degree tail entities and rare relations. These triples have elevated uncertainty despite being in-distribution, reflecting the model’s conservative behavior on under-represented patterns. FP tail entities have mean degree 183 compared to 583 for correctly classified ID triples; FP relations have mean frequency 2,384 vs 5,213.
Appendix G Assumption Verification
Table 12: Empirical verification of theorem assumptions.
@@ Appendix G Assumption Verification
G.1 Empirical Verification of Assumption A3
We empirically verify Assumption A3 (frequency overlap), which is central to Theorem 1’s validity.
@@ G.1 Empirical Verification of Assumption A3
Appendix H OOD Split Methodology
@@ Appendix I Implementation Details
Learned variance alternative. We explored learning relation-conditioned variance $\sigma^{2}(e,r)$ via MLP, avoiding explicit coverage storage. However, this requires auxiliary OOD objectives and provides marginal benefits over explicit coverage. We found CAGP’s interpretability and simplicity preferable.
@@
Table 14: Sensitivity to $\tau$ on FB15k-237 (temporal OOD AUROC).
@@
Appendix J Dataset Statistics
Table 15: Dataset statistics.

### v2

@@ 1 Introduction
Knowledge graph (KG) embeddings enable efficient reasoning over structured knowledge (Bordes et al., 2013; Trouillon et al., 2016). As KGs evolve—new entities emerge, existing entities appear in novel relationships—robust uncertainty quantification becomes critical.
@@
Recent probabilistic methods learn distributions over embeddings (Chen et al., 2019, 2021a, 2021b), where rare entities have higher variance. But these variances are relation-agnostic: entity $e$ receives a single variance $\sigma^{2}_{e}$ regardless of relational context. Consider $e$ appearing in 1000 triples with relations $\{r_{1},\ldots,r_{10}\}$ but never with $r_{11}$. The model assigns low variance (well-observed), yet queries $(?,r_{11},e)$ involve genuinely uncertain predictions.
This causes systematic failures (Figure 1): score-based methods achieve 0.99 AUROC detecting random corruptions (implausible triples) but only 0.52–0.64 on temporal distribution shift (queries from future time periods with new entities and relationships).
@@
Our contributions are threefold: (1) Theoretical: We prove that relation-agnostic uncertainty estimators achieve near-random performance on novel contexts (Theorem 1), and formalize conditions under which semantic and structural uncertainties are complementary (Theorem 2). (2) Empirical: We demonstrate this limitation persists across existing probabilistic methods (Gaussian Process Knowledge Graph Embedding (GP-KGE), Uncertainty-aware Knowledge Graph Embeddings (UKGE), ensembles), achieving only 0.52–0.64 AUROC on temporal shift despite 0.99 on random corruptions. (3) Methodological: We propose CAGP (Coverage-Augmented GP-KGE), which augments entity-level variance with explicit coverage tracking of entity-relation co-occurrence. CAGP achieves 0.94–0.99 AUROC on temporal OOD across multiple benchmarks, a 60–80% relative improvement over probabilistic baselines.
@@ 2 Related Work
UKGE (Chen et al., 2019) associates confidence scores with triples; BEUrRE (Chen et al., 2021a) uses box embeddings where volume indicates uncertainty; GP-KGE (Chen et al., 2021b) learns Gaussian distributions over entities. All share the limitation we identify: learned variances are relation-agnostic. Despite being trained on data containing entity-relation co-occurrence patterns, these methods do not autonomously learn relation-specific uncertainty, necessitating our explicit decomposition framework.
@@
MC Dropout (Gal and Ghahramani, 2016), Deep Ensembles (Lakshminarayanan et al., 2017), and Energy-based methods (Liu et al., 2020b) capture model uncertainty but not structural observation patterns. Simple Neural Gaussian Process (SNGP) (Liu et al., 2020a) estimates distance to training data but remains relation-agnostic. Graph Posterior Networks (Stadler et al., 2021) propagate uncertainty through graphs but treat edges homogeneously.
@@ 3 Background
Given a trained model, distinguish in-distribution (ID) test triples from out-of-distribution (OOD) corruptions. The standard protocol (Safavi and Koutra, 2020) generates OOD triples by replacing the tail with a random entity: $(h,r,t)\to(h,r,t^{\prime})$ where $t^{\prime}\sim\text{Uniform}(\mathcal{E})$. However, this creates easily detectable corruptions (low model scores). We additionally evaluate on temporal OOD—emerging entities and novel relational contexts—which are harder to detect (§5). Performance is measured by AUROC.
@@
Probabilistic embedding methods (Chen et al., 2021b) learn Gaussian distributions over entity embeddings:
@@ 4.1 Preliminaries
We set $\tau$ to the 10th percentile of entity frequencies (robust to outliers; see Appendix C for sensitivity analysis).
@@ 4.3 Theoretical Results
Under assumptions A1–A3 (see Appendix A.1 for full details [^1] We empirically validate Assumption A3 in Appendix A.4.1, finding 100% of novel-context triples have frequency-matched ID counterparts on three benchmarks.), there exists a distribution $\mathcal{D}$ with novel-context OOD subset $\mathcal{D}_{\text{novel}}$ such that:
@@
Under assumptions A1–A6 (see Appendix A.1 for precise statements and robustness analysis):
@@
Proof sketch. Part (i): Novel-context entities have high frequency (by definition), so their variance matches ID entities—semantic uncertainty cannot distinguish them. Part (ii): Emerging entities are defined by low frequency, but may still have coverage for the specific query relation, making some indistinguishable from ID via coverage alone. Part (iii): Novel contexts have missing coverage by definition, while ID has full coverage. Part (iv): The combination uses semantic to catch emerging entities that structural misses (those with coverage), and structural to catch novel contexts that semantic misses. Full proof in Appendix A.1.
4.4 CAGP
@@ 5.1 Experimental Setup
We evaluate on three benchmarks with temporal-like OOD splits: FB15k-237 (14,541 entities, 237 relations), WN18RR (40,943 entities, 11 relations), and YAGO3-10 (123,161 entities, 37 relations). We create temporal-like distribution shifts by partitioning test sets based on entity frequency and entity-relation co-occurrence patterns, simulating the key temporal challenge: detecting when familiar entities appear in unfamiliar relational contexts (methodology detailed in Appendix C.1).
@@
Score-based (UKGE (Chen et al., 2019), Energy (Liu et al., 2020b)), ensemble (MC Dropout (Gal and Ghahramani, 2016), Deep Ensemble (Lakshminarayanan et al., 2017)), distance-aware (SNGP (Liu et al., 2020a)), and single signals ($U_{\text{sem}}$, $U_{\text{str}}$).
@@ 5.4 Standard OOD Detection
Theorem 2 predicts semantic AUROC =0.5=0.5 on novel contexts, but Table 2 shows 0.42. This discrepancy arises because Assumption A3 is approximate: novel-context entities tend toward higher frequencies than the ID average, causing slightly lower semantic uncertainty than ID. The prediction remains directionally correct (near-random).
@@
To characterize CAGP’s failure modes, we analyzed 40,000 predictions on FB15k-237 standard OOD (random tail corruption: 20,000 ID + 20,000 corrupted). The model achieves 0.971 AUROC with 91.3% accuracy. False positives (1,733, 8.7% of ID) occur primarily on triples with low-degree tail entities (mean degree 183 vs 583 for correctly classified ID) and rare relations (mean frequency 2,384 vs 5,213). False negatives (1,733, 8.7% of OOD) involve corrupted tails that coincidentally create higher-degree entities than typical corruptions (mean degree 183 vs 37). The balanced error rates (8.7% FP = FN) and high AUROC validate that both uncertainty components contribute meaningfully. See Appendix B.6 for detailed breakdown.
See Appendix B.1 for additional ablations and Appendix B.3 for calibration analysis.
@@ 6 Conclusion
CAGP’s coverage matrix $\mathbf{C}\in\{0,1\}^{|\mathcal{E}|\times|\mathcal{R}|}$ requires $O(|\mathcal{E}|\times|\mathcal{R}|)$ memory. For our evaluated datasets (FB15k-237: 13MB dense, <<1MB sparse; YAGO3-10: 17.5MB dense, <<1MB sparse), this is negligible. For web-scale KGs, sparse storage (hash tables storing only observed pairs) reduces memory to $O(|\mathcal{T}|)$ with $O(1)$ inference. We also explored learned relation-conditioned variance via multi-layer perceptron, but found explicit coverage simpler and equally effective. CAGP provides interpretability and simplicity suitable for domain-specific KGs with <<1M entities; empirical evaluation on billion-triple KGs remains future work.
@@
Appendix A Theoretical Details
A.1 Theorem Proofs
@@
A.2 Proof of Theorem 1 (Impossibility Result)
@@
A.3 Proof of Theorem 2 (Complementarity)
@@
The assumptions (A1)–(A6) provide idealized conditions ensuring sharp performance guarantees. In practice, these assumptions may be violated to varying degrees. Table 5 (Appendix A.4) quantifies violations across datasets: (A1) Spearman correlation ranges from 0.74-0.74 to 0.85-0.85 (not perfect monotonicity); (A4) bounded semantic gap $\Delta$ exceeds 1.0 on WN18RR ($\Delta=1.10$) and YAGO ($\Delta=1.01$).
@@
A.4 Assumption Verification
Table 5: Empirical verification of theorem assumptions: variance-frequency monotonicity (A1), ID coverage (A2), bounded semantic gap (A4), non-degenerate coverage (A5), and semantic separation (A6).
@@ A.4 Assumption Verification
A.4.1 Empirical Verification of Assumption A3
We empirically verify Assumption A3, which is central to Theorem 1’s validity.
@@ A.4.1 Empirical Verification of Assumption A3
Appendix B Additional Experimental Results
B.1 Ablation Study
Table 7: Ablation of mixing strategies on FB15k-237.
Configuration Temporal OOD Standard OOD
Fixed $\alpha=0.5$ 0.951 0.935
Learned global $\alpha$ (CAGP) 0.986 0.960
B.2 Binary vs Continuous Coverage
Table 8: Binary vs continuous coverage on temporal OOD (FB15k-237). Binary achieves perfect detection (AUROC = 1.0) while continuous variants fail (AUROC = 0.56–0.59).
Coverage Mode AUROC AUPR Separation
Binary 1.0000 1.0000 0.360
Log-scaled 0.5888 0.6119 0.058
TF-IDF 0.5606 0.5733 0.046
Explanation.
Why perfect detection on some temporal splits?
Binary coverage achieves perfect AUROC =1.0=1.0 on FB15k-237’s simulated temporal split (Table 8 above) but only 0.824 on ICEWS14 (main text Table 1). This discrepancy reflects different temporal split compositions (Table 9):
Dataset Novel Contexts Emerging Entities
FB15k-237 (simulated) $\sim$94% $\sim$6%
ICEWS14 (ground-truth) $\sim$61% $\sim$39%
B.3 Calibration Analysis
Method ECE $\downarrow$ Brier $\downarrow$
UKGE 0.312 0.298
Energy 0.287 0.271
Deep Ensemble 0.198 0.203
SNGP 0.183 0.195
$U_{\text{sem}}$ 0.156 0.178
$U_{\text{str}}$ 0.142 0.165
CAGP 0.089 0.112
B.4 Selective Prediction
Method Acc@85% Error Red.
No abstention 62.3%
UKGE 68.1% 15.4%
Deep Ensemble 69.4% 18.8%
$U_{\text{sem}}$ 71.2% 23.6%
$U_{\text{str}}$ 73.5% 29.7%
CAGP 78.4% 42.7%
B.5 Architecture Generalization
Table 12: Generalization across architectures on FB15k-237. CAGP achieves consistent performance (0.959–0.963 AUROC) across DistMult, TransE, and ComplEx.
Base Model $U_{\text{sem}}$ $U_{\text{str}}$ CAGP
DistMult 0.753 0.821 0.959
TransE 0.775 0.822 0.963
ComplEx 0.755 0.821 0.960
B.6 Error Analysis
Metric Value Rate
Overall Performance
AUROC 0.971
Accuracy 91.3%
Precision 0.913
Recall 0.913
F1 Score 0.913
Confusion Matrix
True Negatives 18,267
False Positives 1,733 8.7%
False Negatives 1,733 8.7%
True Positives 18,267
False Positive Characteristics
Avg tail degree 183 (ID: 583)
Avg relation freq 2,384 (ID: 5,213)
False Negative Characteristics
Avg tail degree 183 (OOD: 37)
Key findings.
Table 13 shows that false positives (8.7%) occur primarily on ID triples with low-degree tail entities and rare relations. These triples have elevated uncertainty despite being in-distribution, reflecting the model’s conservative behavior on under-represented patterns. FP tail entities have mean degree 183 compared to 583 for correctly classified ID triples; FP relations have mean frequency 2,384 vs 5,213.
Appendix C Implementation Details
C.1 OOD Split Methodology
@@ C.2 Training Configuration
Learned variance alternative. We explored learning relation-conditioned variance $\sigma^{2}(e,r)$ via multi-layer perceptron, avoiding explicit coverage storage. However, this requires auxiliary OOD objectives and provides marginal benefits over explicit coverage. We found CAGP’s interpretability and simplicity preferable.
@@
Table 14: Sensitivity to $\tau$ (emerging entity threshold, set to 10th percentile of entity frequencies) on FB15k-237 temporal OOD AUROC.
@@
C.3 Dataset Statistics
Table 15: Dataset statistics: number of entities, relations, and triples in train/validation/test splits.